OpenStudy (anonymous):

make r sabject t=r+rV2/8M

OpenStudy (callisto):

t=r+rV^2/8M 8Mt = 8Mr + rV^2 8Mt = r(8M +V^2) r= 8Mt / (8M + V^2)